OpenStudy (anonymous):

A recursive function is shown below: f(1) = 5 and f(n) = f(n - 1) - 6; n > 1 Which of the following lists the terms in the sequence defined by this recursive function? a)5,1,7,13,19 b)5,-1,-7,-13,-19 c)5,11,17,23,29 d)5,-11,-17,-23,-29

OpenStudy (bohotness):

for n=2 you get f(2) = f(2-1) -6 = f(1) -6 = 5-6=-1

OpenStudy (anonymous):

so you just keep plugging in numbers for n ?

OpenStudy (bohotness):

Si

OpenStudy (bohotness):

n=3 f(3) = f(3-1)-6 = f(2) -6 = -1-6=-7

OpenStudy (anonymous):

kk thanks!